# Distributional Scaling Laws for Emergent Capabilities

### v2

@@ Acknowledgments
Appendix A Additional Experimental Details
A.1 Synthetic Tasks
Below we provide more details when training decoder-only language models on the count and reverse order addition—with index hints— tasks. Hyperparameters are largely adapted from (Zhou et al., 2024a). We train all of our models to convergence on the train distribution.
@@ Appendix B Breakthroughness and Linearity
Srivastava et al. (2023) introduced breakthroughness and linearity metrics to capture model performance improving suddenly or reliably with scale. Given a model’s performances $y_{i}$ at model scales $x_{i}$ sorted by ascending model scale, the linearity metric $L$ and breakthroughness metric $B$ are respectively calculated as
@@
Appendix C Additional Figures and Experimental Results
In Figure 10 we sample the five top seeds for the breakthroughness and linearity metric respectively for count (above) and reverse order addition (below).
Our plots in the main paper use a fixed test length 60 for count and fixed test length 40 for reverse order addition. In Figure 11 and Figure 12 we show violin plots to view the random variation distribution at each model scale across multiple test lengths for count and reverse order addition respectively.
@@ Appendix C Additional Figures and Experimental Results
In Figure 8, we saw that for the count task, the bimodal random variation distribution persists when viewing a continuous performance metric: the minimum probability averaged across all test sequences. We show in Figure 14 that a similar outcome holds for the reverse order addition task.
(a) Count.
(b) Reverse order addition.
Figure 10: Top five seeds corresponding to calculations in breakthroughness and linearity given in Appendix B for both the count task (top) and reverse order addition task (bottom).
@@
Figure 11: Violin plots for the same histograms from Figure 5, i.e. for EM accuracy for test length $60$ on the count task, across 8 different scales for independently scaling depth (above) and 10 different scales for scaling width (below).
(a) Fixing network depth to 4 layers.
(b) Fixing network width to 256 hidden dimension.
Figure 12: Violin plots for the same histograms from Figure 2, i.e. for EM accuracy for test length $40$ on the reverse order addition task, across 5 different scales for independently scaling depth (above) and 5 different scales for scaling width (below).
Figure 13: Histogram for EM accuracy for test length 30 (blue, i.e. in-distribution) and test length 60 (orange, i.e. out-of-distribution) on the count task when fixing depth to be one layer. While all model seeds obtain near perfect accuracy in-distribution, all model seeds fail to length generalize at this depth.
(a) Fixing network depth to 6 layers.
(b) Fixing network width to 8 heads per layer.
Figure 14: Histograms of the probabilities assigned to the correct token, calculated by taking the minimum probability within each sequence and averaging across all samples. These histograms are generated for a test length of $40$ on the reverse order addition task task, considering 5 different scaling factors for depth (top) and 3 different scaling factors for width (bottom). As we saw in Figure 8, the bimodal nature of the random variation distribution persists even when measuring probability.
Appendix D Additional LLM Results
@@
Given that the even the performance distributions with the highest variability in Qwen2.5-1.5B are not clearly bimodal, it appears that bimodality depends on model capacity, at least in our continual pretraining setup. We speculate that even a small number of MMLU examples may draw out a latent ability to handle multiple choice formats in the 1.5B model. There is support for the notion that a model might encode some useful latent structure without yet being able to use it; masked language models learn to represent linguistic structure before they learn to use it for complex grammaticality judgments (Chen et al., 2024).

### v3

@@ Acknowledgments
Appendix A Experimental details on synthetic tasks
Below we provide more details when training decoder-only language models on the count and addition—with index hints— tasks. Hyperparameters are largely adapted from [49]. We train all of our models to convergence on the train distribution. Our experiments were run on an internal cluster and all model scales can be run on a single 40GB A100 GPU with gradient accumulation; for count, runs can finish within 2 hours and for addition, runs can finish within 6 hours.
@@ Appendix B Breakthroughness and Linearity
Srivastava et al. [41] introduced breakthroughness and linearity metrics to capture model performance improving suddenly or reliably with scale. Given a model’s performances $y_{i}$ at model scales $x_{i}$ sorted by ascending model scale, the linearity metric $L$ and breakthroughness metric $B$ are respectively calculated as
@@
In Figure 8 we sample the five top seeds for the breakthroughness and linearity metric respectively for count (above) and addition (below).
(a) Count.
(b) Addition.
Figure 8: Top five seeds according to breakthroughness and linearity metrics. Definitions for the two metrics are given in Appendix B, with resulting seeds plotted for 8(a) count and 8(b) addition.
Appendix C Additional Exact Match Accuracy Histogram Plots
(a) Fixing network depth to 4 layers.
(b) Fixing network width to 4 heads per layer.
Figure 9: Random variation in length generalization (count task). Histograms of exact match accuracy on length $60$ sequences when independently scaling 9(a) width and 9(b) depth.
Figure 10: Histograms for EM accuracy when fixing depth to be one layer (count task). We plot the random variation distribution at test length 30 (blue, i.e. in-distribution) and test length 60 (orange, i.e. out-of-distribution). While all model seeds obtain near perfect accuracy in-distribution, all model seeds fail to length generalize at this depth.
Our plots in the main paper fixed test length 40 for addition (Figure 2). In Figure 9 we show analogous histograms for a fixed test length of 60 on the count task. We note that unlike addition, the mode corresponding to failing runs disappears at the highest widths and depths ran, indicating a difference in task difficulty affecting the random variation distribution.
@@ Appendix C Additional Exact Match Accuracy Histogram Plots
Appendix D Additional Continuous Metric Plots
In Section 3.5 we have seen seen that the addition histograms are bimodal in exact match accuracy (Figure 2) and confirmed that the bimodal random variation distribution persists when viewing a continuous performance metric (Figure 6). We can consider another continuous metric, which corresponds to the minimum probability assigned to a token for each sample in the test length generalization dataset:
$\textrm{minprob}(f,X)=\frac{1}{|X|}\sum_{x\in X}\min_{i<|x|}\mathbf{P}_{f}[x_{i}|x_{0\ldots i-1}]$ (2)
We consider the average minimum probability (and analogously average maximum loss) in the sequence because errors are rare, and thus averaging across the sequence obscures generalization failures. However, as the model improves, the lowest-probability token may shift, but this transition is still continuous, not abrupt. Thus, unlike 0/1 accuracy, this metric also avoids threshold effects and better reflects gradual improvements in length generalization, similar to the continuous error metric. We plot the resulting histograms of this metric in Figures 11 and 12 for addition and count respectively. We note that bimodality persists for both tasks, particularly for addition; in particular, stronger bimodality exhibited in the accuracy histograms corresponds to stronger bimodality in the probability histograms.
In Figure 13, we plot the analogous KDE plot of the loss-based error metric for the count task. For count, we note that the mode corresponding to ‘failing runs’ when looking at accuracy or probability is more diffuse compared to addition. However, we claim that stronger bimodality in the original accuracy/probability plots in Figures 9 and 12 is still associated with stronger bimodality in the loss plots. For the addition task, the scales exhibiting the strongest bimodality in accuracy and probability are for width 4 and 8 (in the case of fixed depth) and depth 2 and 4 (in the case of fixed width), and these are also the most strongly bimodal loss KDE distributions. For the count task, even the most strongly bimodal setting (fixed width, with depth 2) has a low probability of failure ( 25%), but still produces a visible wide peak elevated over the long tail of the loss distribution.
(a) Fixing network depth to 6 layers while scaling up width.
(b) Fixing network width to 8 heads per layer while scaling up depth.
Figure 11: Histograms of the minimum probability of any token in each sequence, averaged across sequences (addition task). Metrics are calculated by Equation 2. The bimodal nature of the random variation distribution persists even when using this continuous metric, analogous to Figure 6.
@@ Appendix D Additional Continuous Metric Plots
Figure 12: Histograms of the minimum probability of any token in each sequence, averaged across sequences (count task). Random variation still leads to bimodal performance distributions, even using this continuous performance metric.
(a) Fixing network depth to 4 layers while scaling up width.
(b) Fixing network width to 6 heads per layer while scaling up depth.
Figure 13: Random variation in continuous length generalization error (count task). Kernel Density Estimation (KDE) of loss-based error metric (Equation 1) distribution across model runs. At scales where the EM accuracy distribution is most strongly bimodal, we see that the KDE places the most density at areas of high loss.
Appendix E CSQA LM Results
E.1 Additional Task
In addition to MMLU task, we also reproduce the same data mix experiment on CommonsenseQA (CSQA) [43] task and we include the result in Figure 14. Models yields the same qualitative behaviors as the MMLU task observed in Figure 7.
Figure 14: Violin plots of last-layer reinitialized Qwen2.5 models on the CSQA benchmark. The same behavior observed in MMLU in Figure 7 is reproduced on CSQA as well.
E.2 Impact of Model Size
@@
Given that the even the performance distributions with the highest variability in Qwen2.5-1.5B are not clearly bimodal, it appears that bimodality depends on model capacity, at least in our continual pretraining setup. We speculate that even a small number of MMLU examples may draw out a latent ability to handle multiple choice formats in the 1.5B model. There is support for the notion that a model might encode some useful latent structure without yet being able to use it; masked language models learn to represent linguistic structure before they learn to use it for complex grammaticality judgments [3].